after catching the final out, an outfielder throws the baseball straight up in the air with an initial speed of 35 m/s. How long does it take to reach its highest point?


1
Expert's answer
2021-10-10T15:56:52-0400
"v_y=v_{0y}-gt,""0=v_{0y}-gt,""t=\dfrac{v_{0y}}{g}=\dfrac{35\ \dfrac{m}{s}}{9.8\ \dfrac{m}{s^2}}=3.6\ s."